OpenStudy (asapbleh):

if x=Pi/6. whats 4-2/3 sin (3x-Pi)? I need exact values pls.

OpenStudy (asapbleh):

I got: \[10/3\sin (3\pi/6-\pi)\]

OpenStudy (asapbleh):

But i need help continuing solving it until i get an exact value as a fraction

OpenStudy (asapbleh):

Please help

OpenStudy (anonymous):

\[4-\frac{ 2 }{ 3 }\sin(3x-\pi)=4-\frac{ 2 }{ 3 }\sin(3{\times}\pi/6-\pi)\]\[4-\frac{ 2 }{ 3 }\sin(\pi/2-\pi)=4-\frac{ 2 }{ 3 }\sin(-\pi/2)\]\[4+\frac{ 2 }{ 3 }\sin(\pi/2)\]\[4+\frac{ 2 }{ 3 }\]\[\frac{ 14 }{ 3 }\]

OpenStudy (asapbleh):

THANK YOU VERY MCUH! you just saved my life. I have a test on trig tomorrow. The way you explained it is very throughly. thanks again